The chemical reaction: 4 NH3(g) + 5 O2(g) = 4 NO(g) + 6 H2O(g)is observed to have a rate of reaction with respect to NH3 is 1.5 × 10−4 M / s; what is the rate with respect to H2O? A) −1.5 × 10−4M / s B) +1.5 × 10−4 M / s C) −1.9 × 10−4 M / s D) +1.9 × 10−4 M / s E) +2.3 × 10−4 M / s

A reaction proceeds according to first order kinetics. If the initial concentration is 8.5 M and the rate constant is 0.015 s−1, what is the concentration after 15 seconds?A) 0.09 MB) 0.62 MC) 10.6 MD) 6.8 ME) 82.2 M

Adding 3.00 moles of FeCl3 to 0.50 kg of water results in a solution with a freezing point of −38.0 °C. The water kf = 1.86 °C/m. Solve for the observed van’t Hoff factor for the unknown compound in this solution.A) 2.0B) 2.5C) 3.0D) 3.4E) 30.6
Answer #1

1) The production rate of H2O is calculated:

M H2O = 1.5x10 ^ -4 M NH3 * (6 M H2O / 4 M NH3) = 2.3x10 ^ -4 M

Option E.

2) Using the first order reaction equation, we calculate:

[A] = Exp (-k * t + log [A] o) = Exp (-0.015 * 15 + log 8.5) = 6.8 M

Option D.

3) Molality is calculated:

m = n / Kg = 3 / 0.5 = 6 m

The expected temperature change is calculated:

ΔT = m * Kf = 6 * 1.86 = 11.16 ° C

The factro i is calculated:

i = real ΔT / expected ΔT = 38 / 11.16 = 3.4

Option D.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.